OpenStudy (anonymous):

the derivative of y=x^(2x+1)?

OpenStudy (amistre64):

y=x^(2x+1) Ly= Lx^(2x+1) Ly= (2x+1)Lx 1/y= 2Lx+(2x+1)/x

OpenStudy (anonymous):

the answer is x^(2x+1)*(2lnx+2+1/x) but i can't figure out where the term 2 in the parenthesis came from

OpenStudy (amistre64):

forgot my y' :) 1/y y'= 2Lx+(2x+1)/x y'= y(2Lx+(2x+1)/x)

OpenStudy (anonymous):

ooops

OpenStudy (amistre64):

log it all, the derivae it with logs and product rules

OpenStudy (anonymous):

oh. so by taking the ln of both sides....i can solve it?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

just know that ln(y) ' = 1/y y'

OpenStudy (anonymous):

thank you!!

OpenStudy (amistre64):

yer welcome

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):